# Benign overfitting in leaky ReLU networks with moderate input dimension

### v1

@@ D.1 Training dynamics
D.2 Benign overfitting
Lemma D.1.
Let $\epsilon\in(0,1)$. Suppose that
Then for test data $({\bm{x}},y)$ as per Definition 2.1,
$\mathbb{P}(yf({\bm{W}},{\bm{x}})\leq 0)$ $\leq\epsilon.$
Proof.
Recall that a test point $({\bm{x}},y)$ satisfies
${\bm{x}}=y(\sqrt{\gamma}{\bm{v}}+\sqrt{1-\gamma}{\bm{n}}),$
00 0 $\geq yf({\bm{W}},{\bm{x}})$
$=\sum_{j=1}^{2m}(-1)^{j}y\sigma(\langle{\bm{w}}_{j},{\bm{x}}\rangle)$
When $A_{\min}\geq 0$, this implies that
$\leq 2m(1-\gamma)\sum_{j=1}^{2m}|\langle{\bm{z}}_{j},{\bm{n}}\rangle|^{2}$
$=2m(1-\gamma)\|{\bm{Z}}{\bm{n}}\|^{2}$
$\leq 2m(1-\gamma)\|{\bm{Z}}\|_{F}^{2}\|{\bm{n}}\|^{2},$
where the second inequality is an application of Cauchy-Schwarz. So
By Lemma A.1, the above probability is less than $\epsilon$ if
or equivalently,
Lemma D.2.
$\frac{\ell}{4}\leq|\{i\in\mathcal{S}:y_{i}=1\}|\leq\frac{3\ell}{4}.$
Proof.
$Y=\sum_{i\in\mathcal{I}}Y_{i}=|\{i\in\mathcal{S}:y_{i}=1\}|$
$1-2\exp\left(-\frac{C\ell}{4}\right)\geq 1-\delta$
when $\ell=\Omega\left(\log\frac{1}{\delta}\right)$. ∎
Proof.
$\|{\bm{w}}^{*}\|\leq C\sqrt{\frac{1}{\gamma}+\frac{k}{1-\gamma}}.$
We denote this event by $\omega_{1}$. For $s\in\{1,-1\}$, let $\mathcal{G}_{s}$ denote the set of $i\in\mathcal{G}$ such that $y_{i}=s$. If $n=\Omega\left(\frac{1}{\delta}\right)$ and $k=O(n)$, then $|\mathcal{G}|=\Omega\left(\log\frac{1}{\delta}\right)$. Under these assumptions, by Lemma D.2,
$|\mathcal{G}_{s}|\geq\frac{1}{4}|\mathcal{G}|\geq Cn$
for both $s\in\{1,-1\}$ with probability at least $1-\frac{\delta}{3}$. We denote this event by $\omega_{2}$. For $s\in\{1,-1\}$, let ${\bm{N}}_{s}\in\mathbb{R}^{|\mathcal{G}_{s}|\times d}$ be the matrix whose rows are indexed by $\mathcal{G}_{s}$ and are given by the vectors ${\bm{n}}_{i}$ for $i\in\mathcal{G}_{s}$. As $d=\Omega\left(n\right)=\Omega\left(n+\log\frac{1}{\delta}\right)$ and the rows of ${\bm{N}}_{s}$ are drawn mutually i.i.d. from $\mathcal{N}(\bm{0}_{d},d^{-1}({\bm{I}}_{d}-{\bm{v}}^{T}))$, the following holds by Lemma 4.1. With probability at least $1-\frac{\delta}{3}$ over the randomness of the training data, $\|{\bm{N}}_{s}\|\leq C$ for both $s\in\{1,-1\}$. We denote this event by $\omega_{3}$. Let $\omega=\omega_{1}\cap\omega_{2}\cap\omega_{3}$. By the union bound $\mathbb{P}(\omega)\geq 1-\delta$. We condition on $\omega$ for the remainder of this proof.
$\|{\bm{W}}\|_{F}$ $\leq|\mathcal{A}|\cdot\|{\bm{w}}^{*}\|$
$\leq C|\mathcal{A}|\sqrt{\frac{1}{\gamma}+\frac{k}{1-\gamma}}.$ (15)
11 11 $\leq\hat{y_{i}}f({\bm{W}},{\bm{x}}_{i})$
If $\sqrt{\gamma}A_{\min}\geq\frac{1}{2}$, then
$\frac{A_{\min}}{\|{\bm{Z}}\|_{F}}$ $\geq\frac{A_{\min}}{\|{\bm{W}}\|_{F}}$
$\geq\frac{1}{2\sqrt{\gamma}\|{\bm{W}}\|_{F}}$
$=C\frac{1}{|\mathcal{A}|\sqrt{1+\frac{k\gamma}{1-\gamma}}}$
$\geq C\frac{1}{|\mathcal{A}|+|\mathcal{A}|\sqrt{\frac{k\gamma}{1-\gamma}}}$
$\geq C\frac{1}{|\mathcal{A}|+|\mathcal{A}|\sqrt{\frac{n\gamma}{1-\gamma}}}.$
$\frac{1}{|\mathcal{A}|+|\mathcal{A}|\sqrt{\frac{n\gamma}{1-\gamma}}}\geq C\sqrt{\frac{(1-\gamma)m\log\frac{1}{\epsilon}}{\gamma d}}$
or equivalently
$\sqrt{\frac{(1-\gamma)m\log\frac{1}{\epsilon}}{\gamma d}}+\sqrt{\frac{mn\log\frac{1}{\epsilon}}{d}}$ $\leq\frac{C}{|\mathcal{A}|}.$
This is satisfied when $d=\Omega\left(|\mathcal{A}|^{2}mn\log\frac{1}{\epsilon}\right)$ and $\gamma=\Omega(\frac{1}{n})$, which is true by assumption. So if $\sqrt{\gamma}A_{\min}\geq\frac{1}{2}$, then the network has generalization error less than $\epsilon$ whenever $\omega$ occurs, which happens with probability at least $1-\delta$.
$\leq 2m(1-\gamma)\sum_{j=1}^{2m}|\langle{\bm{z}}_{j},{\bm{n}}_{i}\rangle|^{2}.$
Summing over all $i\in\mathcal{G}_{s}$, we obtain
$=2m(1-\gamma)\sum_{j=1}^{2m}\|{\bm{N}}_{\mathcal{G}_{s}}{\bm{z}}_{j}\|^{2},$
Applying $\omega_{2}$ and $\omega_{3}$, we obtain the bound
$n$ $\leq Cm(1-\gamma)\sum_{j=1}^{2m}\|{\bm{N}}_{\mathcal{G}_{s}}{\bm{z}}_{j}\|^{2}$
$\leq Cm(1-\gamma)\sum_{j=1}^{2m}\|{\bm{z}}_{j}\|^{2}$
$=Cm(1-\gamma)\|{\bm{Z}}\|_{F}^{2}$
$\leq Cm(1-\gamma)\|{\bm{W}}\|_{F}^{2}.$
Then applying (15),
$n$ $\leq Cm(1-\gamma)\|{\bm{W}}\|_{F}^{2}$
$\leq Cm|\mathcal{A}|^{2}\left(\frac{1}{\gamma}+k\right).$
This implies that
$n$ $\leq\frac{Cm|\mathcal{A}|^{2}}{\gamma}$
or
$k$ $\geq\frac{Cn}{m|\mathcal{A}|^{2}}.$
Neither of these conditions can occur if $\gamma=\Omega\left(\frac{m|\mathcal{A}|^{2}}{n}\right)$ and $k=O\left(\frac{n}{|\mathcal{A}|^{2}m}\right)$. Thus, in all cases, the network has generalization error less than $\epsilon$ when $\omega$ occurs, which happens with probability at least $1-\delta$. ∎
D.3 Non-benign overfitting
In this section we show that leaky ReLU networks trained on low-signal data exhibit non-benign overfitting. First we characterize the training dynamics of gradient descent by bounding the noise component learned by the network.
@@ D.3 Non-benign overfitting
As in the case of benign overfitting, we will rely on a generalization bound which depends on the signal-to-noise ratio of the network.
@@ Appendix E Formalizing benign overfitting as a high dimensional phenomenon
Let $(\epsilon,\delta)\in(0,1]^{2}$, $\mathcal{A}$ be a learning algorithm, $\mathcal{M}$ an architecture, $\mathcal{D}$ a data model and $\Omega$ a regime. If for every increasing sequence $(d_{l},n_{l},k_{l},p_{l})_{l\in\operatorname{\mathbb{N}}}\in\Omega^{*}$ there exists an $L\in\operatorname{\mathbb{N}}$ such that for all $l\geq L$ with probability at least $1-\delta$ over $({\bm{X}},\hat{{\bm{y}}})$, where $({\bm{X}},\hat{{\bm{y}}},{\bm{x}},y)\sim D_{d_{l},n_{l},k_{l}}$, it holds that
@@
Let $(\epsilon,\delta)\in(0,1]^{2}$, $\mathcal{A}$ be a learning algorithm, $\mathcal{M}$ an architecture, $\mathcal{D}$ a data model and $\Omega$ a regime. If for every increasing sequence $(d_{l},n_{l},k_{l},p_{l})_{l\in\operatorname{\mathbb{N}}}\in\Omega^{*}$ there exists an $L\in\operatorname{\mathbb{N}}$ such that for all $l\geq L$ with probability at least $1-\delta$ over $({\bm{X}},\hat{{\bm{y}}})$, where $({\bm{X}},\hat{{\bm{y}}},{\bm{x}},y)\sim D_{d_{l},n_{l},k_{l}}$, it holds that

### v2

@@ D.1 Training dynamics
The training dynamics of gradient descent also give us the following result relating the linearization of the noise component of the network to the noise component of the network itself.
@@
D.2 Benign overfitting
Lemma D.2.
Let $\epsilon\in(0,1)$. Suppose that
Then for test data $({\bm{x}},y)$ as per Definition 2.1,
$\mathbb{P}(yf({\bm{W}},{\bm{x}})\leq 0)$ $\leq\epsilon.$
Proof.
Recall that a test point $({\bm{x}},y)$ satisfies
${\bm{x}}=y(\sqrt{\gamma}{\bm{v}}+\sqrt{1-\gamma}{\bm{n}}),$
00 0 $\geq yf({\bm{W}},{\bm{x}})$
$=\sum_{j=1}^{2m}(-1)^{j}y\sigma(\langle{\bm{w}}_{j},{\bm{x}}\rangle)$
When $A_{\min}\geq 0$, this implies that
$\leq 2m(1-\gamma)\sum_{j=1}^{2m}|\langle{\bm{z}}_{j},{\bm{n}}\rangle|^{2}$
$=2m(1-\gamma)\|{\bm{Z}}{\bm{n}}\|^{2}$
$\leq 2m(1-\gamma)\|{\bm{Z}}\|_{F}^{2}\|{\bm{n}}\|^{2},$
where the second inequality is an application of Cauchy-Schwarz. So
By Lemma A.1, the above probability is less than $\epsilon$ if
or equivalently,
Lemma D.3.
$\frac{\ell}{4}\leq|\{i\in\mathcal{S}:y_{i}=1\}|\leq\frac{3\ell}{4}.$
Proof.
$Y=\sum_{i\in\mathcal{I}}Y_{i}=|\{i\in\mathcal{S}:y_{i}=1\}|$
$1-2\exp\left(-\frac{C\ell}{4}\right)\geq 1-\delta$
when $\ell=\Omega\left(\log\frac{1}{\delta}\right)$. ∎
Proof.
$\|{\bm{w}}^{*}\|\leq C\sqrt{\frac{1}{\gamma}+\frac{k}{1-\gamma}}.$
We denote this event by $\omega_{1}$. For $s\in\{1,-1\}$, let $\mathcal{G}_{s}$ denote the set of $i\in\mathcal{G}$ such that $\langle{\bm{v}},{\bm{x}}_{i}\rangle=s$. If $n=\Omega\left(\frac{1}{\delta}\right)$ and $k=O(n)$, then $|\mathcal{G}|=\Omega\left(\log\frac{1}{\delta}\right)$. Under these assumptions, by Lemma D.3,
$|\mathcal{G}_{s}|\geq\frac{1}{4}|\mathcal{G}|\geq Cn$
for both $s\in\{1,-1\}$ with probability at least $1-\frac{\delta}{3}$. We denote this event by $\omega_{2}$. For $s\in\{1,-1\}$, let ${\bm{N}}_{\mathcal{G}_{s}}\in\mathbb{R}^{|\mathcal{G}_{s}|\times d}$ be the matrix whose rows are indexed by $\mathcal{G}_{s}$ and are given by the vectors ${\bm{n}}_{i}$ for $i\in\mathcal{G}_{s}$. As $d=\Omega\left(n\right)=\Omega\left(n+\log\frac{1}{\delta}\right)$ and the rows of ${\bm{N}}_{\mathcal{G}_{s}}$ are drawn mutually i.i.d. from $\mathcal{N}(\bm{0}_{d},d^{-1}({\bm{I}}_{d}-{\bm{v}}^{T}))$, the following holds by Lemma 4.1. With probability at least $1-\frac{\delta}{3}$ over the randomness of the training data, $\|{\bm{N}}_{\mathcal{G}_{s}}\|\leq C$ for both $s\in\{1,-1\}$. We denote this event by $\omega_{3}$. Let $\omega=\omega_{1}\cap\omega_{2}\cap\omega_{3}$. By the union bound $\mathbb{P}(\omega)\geq 1-\delta$. We condition on $\omega$ for the remainder of this proof.
$\|{\bm{W}}\|_{F}$ $\leq|\mathcal{A}|\cdot\|{\bm{w}}^{*}\|$
$\leq C|\mathcal{A}|\sqrt{\frac{1}{\gamma}+\frac{k}{1-\gamma}}.$ (20)
11 11 $\leq\hat{y_{i}}f({\bm{W}},{\bm{x}}_{i})$
If $\sqrt{\gamma}A_{\min}\geq\frac{1}{2}$, then
$\frac{A_{\min}}{\|{\bm{Z}}\|_{F}}$ $\geq\frac{A_{\min}}{\|{\bm{W}}\|_{F}}$
$\geq\frac{1}{2\sqrt{\gamma}\|{\bm{W}}\|_{F}}$
$=C\frac{1}{|\mathcal{A}|\sqrt{1+\frac{k\gamma}{1-\gamma}}}$
$\geq C\frac{1}{|\mathcal{A}|+|\mathcal{A}|\sqrt{\frac{k\gamma}{1-\gamma}}}.$
$\frac{1}{|\mathcal{A}|+|\mathcal{A}|\sqrt{\frac{k\gamma}{1-\gamma}}}\geq C\sqrt{\frac{(1-\gamma)m\log\frac{1}{\epsilon}}{\gamma d}}$
or equivalently
$\sqrt{\frac{(1-\gamma)m\log\frac{1}{\epsilon}}{\gamma d}}+\sqrt{\frac{mk\log\frac{1}{\epsilon}}{d}}$ $\leq\frac{C}{|\mathcal{A}|}.$
This is satisfied for $\epsilon=\exp(-C\cdot\frac{d}{k(1+m|\mathcal{A}|^{2})})$ for some different constant $C$ when $\gamma=\Omega(\frac{1}{k})$, which is true by assumption. So if $\sqrt{\gamma}A_{\min}\geq\frac{1}{2}$, then the network has generalization error less than $\epsilon$ whenever $\omega$ occurs, which happens with probability at least $1-\delta$.
$\leq 2m(1-\gamma)\sum_{j=1}^{2m}|\langle{\bm{z}}_{j},{\bm{n}}_{i}\rangle|^{2}.$
Summing over all $i\in\mathcal{G}_{s}$, we obtain
$=2m(1-\gamma)\sum_{j=1}^{2m}\|{\bm{N}}_{\mathcal{G}_{s}}{\bm{z}}_{j}\|^{2},$
Applying $\omega_{2}$ and $\omega_{3}$, we obtain the bound
$n$ $\leq Cm(1-\gamma)\sum_{j=1}^{2m}\|{\bm{N}}_{\mathcal{G}_{s}}{\bm{z}}_{j}\|^{2}$
$\leq Cm(1-\gamma)\sum_{j=1}^{2m}\|{\bm{z}}_{j}\|^{2}$
$=Cm(1-\gamma)\|{\bm{Z}}\|_{F}^{2}$
$\leq Cm(1-\gamma)\|{\bm{W}}\|_{F}^{2}.$
Then applying (20),
$n$ $\leq Cm(1-\gamma)\|{\bm{W}}\|_{F}^{2}$
$\leq Cm|\mathcal{A}|^{2}\left(\frac{1}{\gamma}+k\right).$
This implies that
$n$ $\leq\frac{Cm|\mathcal{A}|^{2}}{\gamma}$
or
$k$ $\geq\frac{Cn}{m|\mathcal{A}|^{2}}.$
Neither of these conditions can occur if $\gamma=\Omega\left(\frac{1}{k}\right)$ and $k=O\left(\frac{n}{|\mathcal{A}|^{2}m}\right)$. Thus, in all cases, the network has generalization error less than $\exp(-C\cdot\frac{d}{k(1+m|\mathcal{A}|^{2})})$ when $\omega$ occurs, which happens with probability at least $1-\delta$. ∎
We are also able to show the lower bound for the generalization error stated in the main text. See 3.3
Proof.
We proceed along the lines of Theorem 3.2. For $s\in\{1,-1\}$, let $\mathcal{B}_{s}$ denote the set of $i\in\mathcal{B}$ such that $\langle{\bm{v}},{\bm{x}}_{i}\rangle=s$. Note $|\mathcal{B}|=\Omega\left(\log\frac{1}{\delta}\right)$. Under these assumptions, by Lemma D.3,
$|\mathcal{B}_{s}|\geq\frac{1}{4}|\mathcal{B}|\geq Ck$
for both $s\in\{1,-1\}$ with probability at least $1-\frac{\delta}{3}$. We denote this event by $\omega_{1}$. For $s\in\{1,-1\}$, let ${\bm{N}}_{\mathcal{B}_{s}}\in\mathbb{R}^{|\mathcal{B}_{s}|\times d}$ be the matrix whose rows are indexed by $\mathcal{B}_{s}$ and are given by the vectors ${\bm{n}}_{i}$ for $i\in\mathcal{B}_{s}$. As $d=\Omega\left(n\right)=\Omega\left(k+\log\frac{1}{\delta}\right)$ and the rows of ${\bm{N}}_{\mathcal{B}_{s}}$ are drawn mutually i.i.d. from $\mathcal{N}(\bm{0}_{d},d^{-1}({\bm{I}}_{d}-{\bm{v}}^{T}))$, the following holds by Lemma 4.1. With probability at least $1-\frac{\delta}{3}$ over the randomness of the training data, $\|{\bm{N}}_{\mathcal{B}_{s}}\|\leq C$ for both $s\in\{1,-1\}$. We denote this event by $\omega_{2}$. By Lemma D.1, there is a constant $C$ such that
$\|{\bm{Z}}\|_{F}^{2}-2\lambda\sqrt{2m}\|{\bm{Z}}\|_{F}-2m\lambda^{2}$ $\leq\frac{C}{\alpha m}\left(\|{\bm{z}}_{\operatorname{lin}}\|+2m\lambda\right)^{2}.$ (21)
with probability at least $1-\frac{\delta}{3}$. We denote this event by $\omega_{3}$. Let $\omega=\omega_{1}\cap\omega_{2}\cap\omega_{3}$. By the union bound $\mathbb{P}(\omega)\geq 1-\delta$. We condition on $\omega$ for the remainder of this proof.
Let $s\in\{1,-1\}$ be such that $A_{s}=\max\{A_{1},A_{-1}\}$. Since the network attains zero loss, for all $i\in\mathcal{B}_{s}$,
11 11 $\leq\hat{y_{i}}f({\bm{W}},{\bm{x}}_{i})$
$=\sum_{j=1}^{2m}(-1)^{j}y_{i}\sigma(\langle a_{j}{\bm{v}}+{\bm{z}}_{j},-\sqrt{\gamma}y_{i}{\bm{v}}+\sqrt{1-\gamma}{\bm{n}}_{i}\rangle)$
$=\sum_{j=1}^{2m}(-1)^{j}y_{i}\sigma(-\sqrt{\gamma}a_{j}y_{i}+\sqrt{1-\gamma}\langle{\bm{z}}_{j},{\bm{n}}_{i}\rangle)$
$\leq\sum_{j=1}^{2m}(-1)^{j}y_{i}\sigma(-\sqrt{\gamma}a_{j}y_{i})+\sum_{j=1}^{2m}|\sqrt{1-\gamma}\langle{\bm{z}}_{j},{\bm{n}}_{i}\rangle|$
$=\sqrt{\gamma}\sum_{j=1}^{2m}(-1)^{j+1}s\sigma(sa_{j})+\sqrt{1-\gamma}\sum_{j=1}^{2m}|\langle{\bm{z}}_{j},{\bm{n}}_{i}\rangle|$
$=-\sqrt{\gamma}A_{s}+\sqrt{1-\gamma}\sum_{j=1}^{2m}|\langle{\bm{z}}_{j},{\bm{n}}_{i}\rangle|.$
From which we conclude
$\sqrt{1-\gamma}\sum_{j=1}^{2m}|\langle{\bm{z}}_{j},{\bm{n}}_{i}\rangle|\geq 1+\sqrt{\gamma}A_{s}\geq\sqrt{\gamma}A_{s}$
for all such $i$. Squaring both sides of the inequality and applying Cauchy-Schwarz, we obtain
$\gamma A_{s}$ $\leq(1-\gamma)\left(\sum_{j=1}^{2m}|\langle{\bm{z}}_{j},{\bm{n}}_{i}\rangle|\right)^{2}$
$\leq 2m(1-\gamma)\sum_{j=1}^{2m}|\langle{\bm{z}}_{j},{\bm{n}}_{i}\rangle|^{2}.$
Summing over all $i\in\mathcal{B}_{s}$, we obtain
$|\mathcal{B}_{s}|\gamma A_{s}$ $\leq 2m(1-\gamma)\sum_{i\in\mathcal{B}_{s}}\sum_{j=1}^{2m}|\langle{\bm{z}}_{j},{\bm{n}}_{i}\rangle|^{2}$
$=2m(1-\gamma)\sum_{j=1}^{2m}\|{\bm{N}}_{\mathcal{B}_{s}}{\bm{z}}_{j}\|^{2},$
Applying $\omega_{2}$ and $\omega_{3}$, we obtain the bound
$k\gamma A_{s}$ $\leq Cm(1-\gamma)\sum_{j=1}^{2m}\|{\bm{N}}_{\mathcal{B}_{s}}{\bm{z}}_{j}\|^{2}$
$\leq Cm(1-\gamma)\sum_{j=1}^{2m}\|{\bm{N}}_{\mathcal{B}_{s}}\|^{2}\|{\bm{z}}_{j}\|^{2}$
$\leq Cm(1-\gamma)\sum_{j=1}^{2m}\|{\bm{z}}_{j}\|^{2}$
$=Cm(1-\gamma)\|{\bm{Z}}\|_{F}^{2}.$
For $k=\Omega(\frac{1}{\alpha})$, this inequality along with Assumption 1 implies that
$C\|{\bm{Z}}\|_{F}^{2}\leq\|{\bm{Z}}\|_{F}^{2}+2\lambda\sqrt{2m}\|{\bm{Z}}\|_{F}+2m\lambda^{2}$
for a different constant $C$. With the last two inequalities and equation 21, we obtain the bound, for a new constant $C$.
$k\gamma A_{s}$ $\leq C\frac{1-\gamma}{\alpha}\left(\|{\bm{z}}_{\operatorname{lin}}\|+2m\lambda\right)^{2}.$
We then apply $k=\Omega(\frac{1}{\alpha})$ and Assumption 1 again to conclude that for some $C$,
$\|{\bm{z}}_{\operatorname{lin}}\|\geq C\sqrt{\frac{k\gamma A_{s}\alpha}{1-\gamma}}.$
Note that
$A_{\operatorname{lin}}=\frac{A_{1}+A_{-1}}{1+\alpha}\leq 2A_{s}.$
We then bound
$\frac{A_{\operatorname{lin}}}{\|{\bm{z}}_{\operatorname{lin}}\|}\leq C\frac{A_{s}}{\sqrt{\frac{k\gamma A_{s}\alpha}{1-\gamma}}}$
$\leq C\sqrt{\frac{1-\gamma}{k\gamma\alpha}}$
for some constant $C$. Now consider a test point $({\bm{x}},y)$, which satisfies
${\bm{x}}=y(\sqrt{\gamma}{\bm{v}}+\sqrt{1-\gamma}{\bm{n}}),$
where ${\bm{n}}\sim\mathcal{N}(\textbf{0}_{d},\frac{1}{d}({\bm{I}}_{d}-{\bm{v}}{\bm{v}}^{T}))$. Since the data distribution is symmetric,
$\mathbb{P}(yf({\bm{W}},{\bm{x}})\leq 0)$ $\geq\frac{1}{2}\mathbb{P}(yf({\bm{W}},{\bm{x}})\leq 0\text{ or }-yf({\bm{W}},-{\bm{x}})\leq 0)$
$\geq\frac{1}{2}\mathbb{P}(yf({\bm{W}},{\bm{x}})-yf({\bm{W}},-{\bm{x}})\leq 0).$
We see that
$yf({\bm{W}},{\bm{x}})-yf({\bm{W}},-{\bm{x}})=(1+\alpha)\left(yA_{\operatorname{lin}}\sqrt{\gamma}+\langle{\bm{z}}_{\operatorname{lin}},{\bm{n}}\rangle\sqrt{1-\gamma}\right)$
By Lemma C.1 we then can bound
$\mathbb{P}(y\langle{\bm{w}},{\bm{x}}\rangle\leq 0\;|\;\omega)$ $\geq\frac{1}{8}\exp\left(-\frac{6d}{\pi}\frac{\gamma}{1-\gamma}\frac{A_{\operatorname{lin}}^{2}}{\|{\bm{z}}_{\operatorname{lin}}\|^{2}}\right)$
$\geq\exp\left(-\frac{Cd}{\alpha k}\right)$
for a new constant $C$, provided $A_{\operatorname{lin}}$ is positive. In the last line we can bound $\frac{d}{k}$ below as $d=\Omega(n)=\Omega(k)$. If $A_{\operatorname{lin}}$ is negative, then the generalization error is at least $\frac{1}{4}$ which is also at least $\exp(-Cd/(\alpha k))$. ∎
D.3 Non-benign overfitting
@@
In this section we show that leaky ReLU networks trained on low-signal data exhibit non-benign overfitting. As in the case of benign overfitting, we will rely on a generalization bound which depends on the signal-to-noise ratio of the network.
@@ Appendix E Formalizing benign overfitting as a high dimensional phenomenon
Let $(\epsilon,\delta)\in(0,1]^{2}$, $\mathcal{A}$ be a learning algorithm, $\mathcal{M}$ an architecture, $\mathcal{D}$ a data model and $\Omega$ a regime. If for every increasing sequence $(d_{l},n_{l},k_{l},p_{l})_{l\in\operatorname{\mathbb{N}}}\in\Omega^{*}$ there exists an $L\in\operatorname{\mathbb{N}}$ such that for all $l\geq L$ with probability at least $1-\delta$ over $({\bm{X}},\hat{{\bm{y}}})$, where $({\bm{X}},\hat{{\bm{y}}},{\bm{x}},y)\sim D_{d_{l},n_{l},k_{l}}$,
it holds that
@@
Let $(\epsilon,\delta)\in(0,1]^{2}$, $\mathcal{A}$ be a learning algorithm, $\mathcal{M}$ an architecture, $\mathcal{D}$ a data model and $\Omega$ a regime. If for every increasing sequence $(d_{l},n_{l},k_{l},p_{l})_{l\in\operatorname{\mathbb{N}}}\in\Omega^{*}$ there exists an $L\in\operatorname{\mathbb{N}}$ such that for all $l\geq L$ with probability at least $1-\delta$ over $({\bm{X}},\hat{{\bm{y}}})$, where $({\bm{X}},\hat{{\bm{y}}},{\bm{x}},y)\sim D_{d_{l},n_{l},k_{l}}$,
it holds that